OpenStudy (anonymous):

-5x/3x+3=2x/b+6+6x-4/x+1

OpenStudy (anonymous):

is that really a "b" or is it an "x"

OpenStudy (anonymous):

\[\frac{-5}{3x+3}=\frac{2x}{x+6}-\frac{4}{x-1}\]?